OpenStudy (anonymous):

The width of a rectangular plot is 4 feet longer than two-fifths the length. If the perimeter of the plot is 36 feet, find the length

OpenStudy (anonymous):

make some expressions width=4 feet longer than 2/5 of the length so width=(2/5)L+4 length=L perimeter=2L+2W plug in our expression for w 2L+2((2/5)L+4) can you solve for L?

OpenStudy (anonymous):

I'm seriously really confused

OpenStudy (anonymous):

!!!!!